TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.AntonioVergara on 2007-06-27 - 19:09

Powering Tests Sector 7-8 - Wednesday 27.06.2007

PDF Progress_PoweringS78_070628.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x06x27x18x48/Progress_PoweringS78_070628.pdf

Test Report from Wednesday 27.06.2007

  • RB.A78
    • PLI1.4 - FPA from PIC - Test OK
    • PLI1.5 - Quench heater discharge on A28.R7 - Test OK
    • The bus-bar detection threshold is increased from 1V to 2V in order to ease the ramp up
    • Circuit successfully ramped to 2000A!!
      • PLI2.1 - Current cycle - Test OK
      • PLI2.2 - Energy extraction discharge (quench to a current lead) - Test OK
      • PLI2.2b - SPA from PIC - Test OK
      • PLI2.3 - Quench heater discharge on LBBRA.13L8 (magnet 3002) and LBBRA.13R7 (actually we wanted to quench LBARA.13R7) - Test OK

  • RQF.A78 and RQD.A78 ramped up together to 6500A. Quench on 15L8 - Test OK

  • Simmultaneous powering of RQF.A78, RQD.A78 and RB.A78 - Strated @ 21:00 - Stopped @ 22:00 - Test OK

  • 600A Correctors - 9 new circuits released by QPS (see pdf file attached)
    • PIC2 done on 9 circuits
    • PCS.2/.3/.4 on RQS.L8B1 and RCO.A78B1

  • 60A Closed Orbit Correctors
    • P2N continued for 2 circuits

  • Matching Section
    • Reproduction of the quench event of May 31st - Magnet did not quench.
    • Last commissioning step of RQ5.L8 via the sequencer - Test OK

  • Stability test on RQ4.L8B1/B2

Test Plan for Thursday 28.06.2007 - Noon Update

Tests will resume once ElQA and connection of the current leads on 8 circuits of the DFBAO (see file attached) are completed.

  • PCS tests for all the circuits released by PIC2

  • In parallel: PIC2 for the eight circuits that have not gone through it yet (whenever possible)

  • RQF.A78, RB.A78 - Performance tests

  • Local tests of AB/PO in one corrector circuit of DFBMC (the test implies disconnection-connection of active parts in the converter side)

  • Matching Section: Powering of RD2.L8 and RQ4.L8 with DFBMA at minimum liquid operational level (31%) - RD2.L8 quench detector level at 20mV

  • P2N continues for 60 closed orbit correctors

  • Quadruple quadrupole quench post-poned for tomorrow before lunch.

Outstanding activities

  • Performance tests of two 600A circuits once released, 4 hours per circuit are required.
  • Tracking tests of RB/QF/QD using PO function generation. Provisionally scheduled for Friday 29/07, from 14:00
  • Tracking tests of all available circuits (RB/QF/QD/Q4/Q5/D2/600A) using LSA application. Provisionally scheduled for Monday 2/7/07, from 08:00
  • Long duration tests of all available electrical circuits, minimum of 8H test
  • ElQA for circuits not powered in this run
  • AUG - Scheduled for Monday 9/7/07

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-06-27 - 18:52

going on activities in sector 45

Going on activities in sector 45

  • DFBAH and MB (diode)
    • checked and closed by Wednesday 27th;
    • Leak test on Thursday 28th;
    • Purging on Friday 29th evening and all the week end long;
    • Refilling on Monday 2nd July;
    • On Tuesday 3th July, once 6Bar conditions are reached, ElQA will be repeated (only on RB);
  • CRYO-instrumentation
    • Initialization of RF valves on Wednesday 27th or Thursday 28th;
    • Coherence test of all the valves in ARC and LSS (Thursday 28th and Friday 29th);
    • Commissioning;
    • Initialization of the CV910 valves on Tuesday 26th and Wednesday 27th;
  • TS/EL intervention
    • Changing of the connectors of a BPM (LSS4R);
    • Placing of some Radmon cables on Wednesday 27th, Thursday 28th and Friday 29th (LSS5L);
    • Connection check of the instrumentation cables of BCT;
  • TIM camera
    • General check on Monday 2nd July;
    • Run from P4 to P5, during the Cryo Patrouille the morning of the start of the cool-down;
  • Safety Inspection
    • Monday 2nd July at 2pm;
  • RAMSES
    • Check of the equipment on Tuesday 26th and Wednesday 27th.
  • CV Intervention
    • Thursday 28th June from 2pm to 4pm
  • QUI
    • Caps removal, sleeve installation and then closure of the QUI performed on Thursday 28th.
    • The new mechanism for the header B valve is being prepared and should be ready for Friday 29th or early next week.

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login